### v2

# Generate, but Verify: Reducing Hallucination in Vision-Language Models with Retrospective Resampling

Abstract
Vision-Language Models (VLMs) excel at visual understanding but often suffer from visual hallucinations, where they generate descriptions of nonexistent objects, actions, or concepts, posing significant risks in safety-critical applications. Existing hallucination mitigation methods typically follow one of two paradigms: generation adjustment, which modifies decoding behavior to align text with visual inputs, and post-hoc verification, where external models assess and correct outputs. While effective, generation adjustment methods often rely on heuristics and lack correction mechanisms, while post-hoc verification is complicated, typically requiring multiple models and tending to reject outputs rather than refine them. In this work, we introduce REVERSE, a unified framework that integrates hallucination-aware training with on-the-fly self-verification. By leveraging a new hallucination-verification dataset containing over 1.3M semi-synthetic samples, along with a novel inference-time retrospective resampling technique, our approach enables VLMs to both detect hallucinations during generation and dynamically revise those hallucinations. Our evaluations show that REVERSE achieves state-of-the-art hallucination reduction, outperforming the best existing methods by up to 12% on CHAIR-MSCOCO and 34% on HaloQuest.
[…]
1 Introduction
Vision-Language Models (VLMs) have revolutionized visual understanding, achieving dramatic improvements in tasks like visual question-answering and image captioning, yet they still struggle with a significant limitation: visual hallucination — the tendency to describe objects that aren’t actually present in the scene. Such hallucinations pose significant risks when applying VLMs to safety-critical environments, ranging from autonomous driving scenarios and decision-making to assistive technologies for the visually impaired.
To tackle these issues, researchers have generally pursued methods following one of two paradigms: generation adjustment or post-hoc verification. Generation adjustment methods focus on aligning textual outputs more closely with visual inputs by modifying the VLM’s generation behavior, either in a “training-free” way (modifying the logits at decoding time) [27, 22, 23, 3, 56], or using a “training-based” strategy requiring additional supervision or custom objective functions [52, 40, 54, 39, 32, 53]. Unfortunately, these methods have no means of correcting erroneous tokens once they have been generated, and they do not leverage powerful retrospective tools such as chain-of-thought reasoning to reason about and evaluate the quality of their generation. In contrast to generation adjustment approaches, post-hoc verification methods [51, 55, 36, 47, 40] leverage large external models, such as GPT-4 [35], as verifiers to evaluate outputs after they have been generated. Post-hoc verifiers are accurate at predicting hallucination, but are complicated, requiring multiple models. Post-hoc verifiers often do not provide a way for the model to correct the hallucination but instead adopt generic refusal strategies.
In this paper, we introduce REVERSE (REtrospective VERification and SElf-correction), the first framework that unifies generation adjustment with online post-hoc verification in a single VLM architecture (Figure 1). REVERSE consists of two key components. First, our method fine-tunes a VLM on a specially constructed training dataset consisting of synthetic hallucination phrases tagged by a special token. Unlike prior VLMs instructed with only well-grounded data, our resulting hallucination-aware model is now able to tag likely phrase-level hallucinations during the generation process.
Second, we introduce retrospective resampling, a technique that allows the hallucination-aware VLM to serve as its own verifier. During the generation process, when the hallucination-aware VLM places sufficient probability on the special hallucination token we trigger a backtracking self-correction process. Specifically, we backtrack to a previous confident section and then apply rejection sampling and query-rewriting to correct the hallucination. As illustrated in Figure 1, both components in REVERSE combine to effectively mitigate visual hallucinations.
We evaluate REVERSE against SOTA hallucination reduction baselines across a wide range of benchmarks designed for hallucination evaluation on LLaVA-v1.5 [33], LLaVA-MORE [16], and Qwen2.5-VL [6]. On captioning tasks, REVERSE achieves up to a 12% reduction in CHAIR scores on CHAIR-MSCOCO [38] and AMBER [45] over the best existing methods. On hallucination-sensitive open-ended tasks, it also delivers over a 10% and 34% performance improvement on MMHal [40] and HaloQuest [47], respectively.
In summary, this paper both (i) introduces REVERSE, the first hallucination reduction method unifying the generation adjustment and post-hoc verification approaches, addressing hallucination in both the training and inference stages and (ii) provides a new public training dataset and data curation pipeline for training-time hallucination mitigation consisting of 1.3M semi-synthetic samples. Together, these contributions allow REVERSE to achieve up to a 12% improvement on the CHAIR-MSCOCO benchmark, and a 34% improvement on the HaloQuest benchmark over existing SOTA methods for hallucination reduction under the same setting, especially in questions with false premise and insufficient context.
2 Background & Related Work
Following the success of Large Language Models (LLMs) [42, 4, 5, 31, 11, 1], Vision-Language Models (VLMs) have shown success across various multimodal tasks, such as image captioning, visual question answering, visual reasoning, and image segmentation [28, 2, 33, 34, 35, 41, 6, 50, 49, 26, 46]. Despite their impressive performance, VLMs are prone to hallucinations: generating incorrect or nonexistent visual information [29]. To address this issue, several hallucination-specific benchmarks, such as CHAIR-MSCOCO [38], AMBER [45], MMHal [40], and POPE [29], and HaloQuest [47] have been introduced. These benchmarks evaluate VLM hallucinations across both discriminative and generative tasks, with a growing trend of using VLMs for automatic visual hallucination detection.
Beyond detection, several recent methods attempt to mitigate hallucinations by adjusting a VLM’s generation process. Training-free approaches primarily focus on improving decoding strategies [27, 22, 23, 3, 56]. For instance, VCD [27] employs contrastive decoding, OPERA [22] introduces a penalty term during beam search, and DoLA [15] enhances decoding by contrasting different model layers. Training-based methods, on the other hand, aim to reduce hallucinations through improved training objectives and additional data. Some approaches leverage data augmentation [10], while others refine training via reinforcement learning from human feedback (RLHF) [40, 52, 54]. Additional methods fine-tune VLMs using custom loss functions, such as EOS token-based penalties for lengthy descriptions [53], contrastive learning from paired correct-hallucination data [23], and visual instruction tuning with improved datasets [32]. However, these approaches merely adjust the generator’s behavior rather than fundamentally eliminating hallucinations—once incorrect information is produced, there is no built-in mechanism for correction.
The closest prior work to ours includes Woodpecker [51] and LURE [55], which use external models to verify and rewrite initial outputs from a VLM. While effective to some extent, these methods rely on complex, multi-stage pipelines with external dependencies. Moreover, they suffer from error propagation, as a single-round rewriting step is often insufficient to fully recover from low-quality initial outputs.
In contrast, our method is the first unified framework where the VLM itself serves as both the generator and verifier, enabling self-correction in a streamlined and integrated manner. Compared to prior generation-adjustment methods, our self-verification pipeline allows VLMs to retrospect and iteratively self-correct after content has been generated. Compared to existing post-hoc refinement approaches, our method eliminates the need for external models or complex multi-stage pipelines, achieving better results as the verifier can instantly and iteratively correct the generator’s outputs.
3 REVERSE: Retrospective Verification and Self-Correction
REVERSE (REtrospective VERification and SElf-correction) is a hallucination reduction paradigm for Vision-Language Models (VLMs) that unifies generation adjustment and post-hoc verification methods. REVERSE allows VLMs to be hallucination-aware by explicitly modeling and monitoring the likelihood that each generated phrase is well-grounded. During training, the model is explicitly trained to classify each groundable phrase as either “confident” or “unconfident” and during inference, the model generates responses while continuously verifying the confidence of each phrase using the likelihood of the “unconfident” predictor. If a phrase is sufficiently ungrounded, the model then performs retrospective adjustment to refine the segment, enabling self-correction on the fly.
Key to the first goal of classifying each phrase as “confident” or “unconfident” is training the model to understand if a phrase is well-grounded. While VLMs and LLMs inherently provide implicit confidence scores through token probabilities, these scores are often mis-calibrated and do not consistently correlate with output correctness, making them unreliable for verification [48, 14, 17]. Furthermore, even when accurate, these probabilities offer no indication of where to backtrack for phrase re-generation and self-correction.
To overcome these limitations, we introduce three tokens to the VLM vocabulary that can be used to explicitly mark key phrases and represent the model’s confidence level:
<SPAN>: Marks the beginning of key or object phrases.
</CN>: Marks the end of confident, grounded phrases.
</UN>: Marks the end of unconfident, hallucinated phrases.
These tokens, when placed before/after objects or phrases in the scene can serve as ad-hoc classifiers of the confidence of the model. I.e. if a model generates a </UN> token after a phrase, that phrase can be considered to be ungrounded, while if it generates a </CN>, that phrase is likely grounded in the image. Annotating our data with such tokens, as is shown in Figure 2, will allow us to train the VLM itself to perform post-hoc verification instead of relying on an external model.
3.1 Data Curation
Towards models that are capable of automatically tagging phrases as confident or unconfident, we constructed a 1.3M VLM instruction-tuning dataset containing a total of 6.8M question-answer pairs (turns), of which 3.8M are correct answers and 2.9M are hallucinated answers. In each response, all positive phrases are enclosed with <SPAN> and </CN> while the negative phrases are enclosed with </UN>.
The phrases in this dataset are automatically generated by annotating existing training data for VLM with these tokens. Our dataset is initially sourced from the LLaVA-v1.5-665k instruction tuning data [34], which contains only “positive” or “well-grounded” samples. To introduce “negative” or “un-grounded” samples, we designed a multi-step pipeline that generates incorrect answers leveraging rule-based algorithms and gpt-4o-mini-0718 [35]. Specifically, we first classify the answer types, in which we can augment most of the answers with rule-based methods easily, such as binary Yes/No questions or counting questions. For the remaining general answers, mostly long answers or descriptions, we apply AI inference for high-quality and diverse data augmentation. For negative samples, constrain the sentence to immediately terminate upon reaching </UN>. This design not only prevents VLMs from continuing to generate ungrounded descriptions after detecting hallucinated content but also helps maintain training data quality, as any remaining context may become meaningless once the preceding information has been altered. To support retrospective query re-writing, we further inject negative keywords as hints (further discussed in subsection 3.3). Our dataset is twice the size of the LLaVA-v1.5-665k instruction tuning dataset while maintaining a similar overall composition. It preserves the same average question-answer pairs per sample and a comparable question type distribution. More details about the dataset/dataset generation pipeline are provided in Appendix B.
3.2 Hallucination-aware Training
To train REVERSE to recognize and respond to the new tokens, we introduce a modified cross-entropy next-token prediction loss that prevents hallucination while modeling confidence levels. Our training objectives are threefold. First, we aim to enable conventional instruction tuning to allow VLMs to perform next-token prediction to generate accurate answers. Second, we wish to reduce the likelihood of generating the hallucinated tokens that we have introduced in the new dataset. Third, we want to teach the model to generate <SPAN> at the start of key phrases, and </CN> or </UN> as explicit confidence estimators around those phrases. We achieve all of these goals by assigning a weight to each token during training; positive weights are assigned to tokens outside the <SPAN>… </UN> bounds, encouraging standard next-token prediction while zero-weights are assigned to tokens within <SPAN> and </UN> (i.e., masking out the targets) to avoid impacting the likelihood when training on ungrounded data (and reinforcing language priors).
Formally, let be our model and be the labeled VQA dataset, where each sample consists of an input sequence and an output sequence . Here, includes both encoded image features and question (query) tokens, while each in can be either a text token corresponding to the answer or one of the three special tokens.
The model predicts the next token probability as . We then define the modified negative log-likelihood loss for a given sample as:
where is an indicator variable taking the value for all tokens except those enclosed by <SPAN> and </UN>. For tokens within these markers, . Note that this masking is applied to the targets rather than the inputs. We optimize model parameters using the loss in Equation 1. The full training procedure is described in section 4.
3.3 Retrospective Resampling
During inference, the model follows standard next-token prediction but continuously monitors the likelihood of </UN>, triggering retrospective resampling when that likelihood exceeds a pre-defined threshold (see Figure 3). Specifically, whenever </CN> or </UN> is generated (the end of a span), we compute the probability of </UN>, denoted as , across previous tokens. If surpasses a predefined threshold , the model initiates a self-correction process via backtracking and retrospective resampling. Otherwise, generation proceeds normally, with <SPAN>, </CN> and </UN> tokens removed before presenting the final output.
Backtracking Strategies
A critical challenge in self-correction is determining both (1) where to backtrack and (2) how to regenerate content. To determine where to backtrack to, our approach follows a hierarchical fallback strategy:
1.
The model first backtracks to the most recent </CN>, which attempts to adjust only the local information to reduce the likelihood of hallucination.
2.
If the issue persists after local correction attempts, it is likely that the hallucination issue stems from earlier information in the sequence. Thus, we revert further, backtracking to the last sentence boundary (indicated by the last punctuation token).
3.
If self-correction continues to fail after total attempts, the output is finalized and returned to the user, along with an indication that a hallucination was detected, but could not be corrected.
Since is typically low (even under hallucinations), explicitly waiting for </UN> to appear in the output is impractical. Instead, we set a confidence threshold , which allows proactive identification of hallucinated phrases before they fully form. The effect of selection is analyzed in subsection 4.2. After the backtracking has occurred, we leverage rejection sampling and query rewriting for self-correction.
Rejection Sampling
Rejection sampling refines uncertain phrases by resampling multiple times at an increased temperature, seeking an alternative where remains below . The process continues until reaching a confident phrase (marked by </CN>) or exhausting the maximum resampling attempts. In this work, we make no attempts to “ban” the generation of the same tokens during the re-sampling process. This procedure, while potentially more efficient, often leads to issues where innocuous tokens such as “a” or “the” are banned, leading to disfluencies in the final generated text. Instead, we rely on the increased temperature to lead to new candidates over several repeated generations. While rejection sampling is effective for resolving localized hallucinations, its success depends on the ability of the model to generate valid alternatives within a reasonable number of attempts. In cases where repeated resampling fails to produce an acceptable phrase, the model may need to fall back on broader correction strategies, such as query rewriting, to address deeper inconsistencies in the generated content.
Query Rewriting
In addition to rejection sampling, we found that “query rewriting” can provide stronger signals for VLMs to do self-correction. Query rewriting dynamically modifies the prompt to encourage better factual grounding. Specifically, the input prompt is augmented with a clarification hint:
<system-prompt> [<optional image>] <question> (Hint: potential incorrect phrases → <placeholder>)
This prompt signals the model to reconsider flagged segments and generate a more reliable response. In addition to rejection sampling with increased temperature, which iteratively refines outputs by resampling under varied decoding conditions, query rewriting can directly influence the model’s contextual understanding by reformulating its input conditions. Since our training data includes hallucination-corrected phrase pairs, we randomly inject 20% of this query-rewriting prompt into the instruction-tuning process. This improves the model’s ability to recognize the hint, making retrospective resampling more effective.
| Base VLM | Method Type | Method | CHAIR-MSCOCO | AMBER-G | ||||
| CHAIR | Cover | Hall | Cog | |||||
| LLaVA-v1.5 7B [34] | None | 15.4 | 50.0 | 7.8 | 51.0 | 36.4 | 4.2 | |
| Gen-Adjust | VCD [27] | 14.9 | 48.6 | - | - | - | - | |
| OPERA‡ [22] | 14.6 | 47.8 | 7.3 | 49.6 | 32.0 | 3.5 | ||
| DoLA [15] | 14.1 | 51.6 | 7.6 | 51.6 | 36.0 | 4.0 | ||
| AGLA [3] | 14.1 | 43.0 | - | - | - | - | ||
| MEMVR [56] | 13.0 | 46.6 | - | - | - | - | ||
| w/ Train | EOS [53] | 12.3 | 40.2 | 5.1 | 49.1 | 22.7 | 2.0 | |
| HALVA [39] | 11.7 | 41.4 | 6.6 | 53.0 | 32.2 | 3.4 | ||
| HA-DPO [54] | 11.0 | 38.2 | 6.7 | 49.8 | 30.9 | 3.3 | ||
| Post-hoc Refine | Woodpecker† [51] | 14.8 | 45.8 | 6.9 | 48.9 | 30.4 | 3.6 | |
| Combination | REVERSE() | 10.3 | 37.0 | 6.0 | 52.2 | 30.4 | 3.0 | |
| REVERSE() | 6.1 | 13.6 | 4.0 | 26.9 | 10.2 | 0.9 | ||
| LLaVA-MORE 8B [16] | None | 14.4 | 52.0 | 7.8 | 53.1 | 36.6 | 3.9 | |
| DoLA [15] | 13.8 | 51.8 | 7.9 | 53.1 | 38.4 | 4.1 | ||
| Woodpecker [51] | 14.3 | 51.0 | 7.4 | 50.7 | 36.7 | 3.7 | ||
| REVERSE() | 12.2 | 42.4 | 6.5 | 54.8 | 35.5 | 3.9 | ||
| REVERSE() | 8.4 | 25.2 | 5.1 | 38.9 | 20.8 | 2.1 | ||
| Qwen2.5-VLFT 3B [6] | None | 12.2 | 45.8 | 7.7 | 51.7 | 35.9 | 4.1 | |
| DoLA [15] | 14.0 | 47.6 | 9.7 | 48.1 | 31.4 | 1.9 | ||
| REVERSE() | 10.5 | 39.4 | 7.5 | 51.5 | 34.4 | 3.6 | ||
4 Experiments
Implementation Details We applied our method, REVERSE, on three VLM backbones: LLaVA-v1.5 (7B) [34], LLaVA-More (LLaVA with Llama-v3.1 8B) [16, 1], and Qwen2.5-VL (3B) [6]. Since LLaVA provides both its pre-trained model and instruction tuning data, we performed LoRA fine-tuning on the pre-trained model directly with our modified cross-entropy loss (see subsection 3.2) and the 1.3M-sample dataset for one epoch. In contrast, Qwen2.5-VL does not release its instruction tuning data. To enable a fair comparison, we perform full fine-tuning on the publicly available Qwen2.5-VL model using two alternatives: a 100k subset of LLaVA’s instruction data and a matched subset from our dataset. Although a more direct evaluation using Qwen2.5-VL’s original instruction tuning data and the augmentation method described in subsection 3.1 would be ideal, it is not feasible given the current release conditions. More details on training recipes are provided in Appendix F.
During inference, we apply retrospective resampling with different threshold values: for LLaVA-series models and for Qwen2.5-VL. These values are set per model backbone, as confidence scores across LLMs and VLMs are typically not calibrated and are rarely shared due to differences in training [24, 12]. To ensure fairness, we apply a consistent threshold per model across all evaluation datasets. Further discussion on how this controllable threshold affects model behavior is provided in subsection 4.2. For the correction mechanism, we allow up to a total correction attempts, with local correction attempts of . Additionally, we implement rejection sampling with a base temperature of , gradually increasing it with a step size of , capped at a maximum temperature of : .
Evaluation Protocol To evaluate our approach, we compare REVERSE with various hallucination mitigation methods, including training-free or training-based generative adjustment techniques [27, 22, 15, 3, 53, 39, 54, 56], and post-hoc verification with refinement [51]. All methods are evaluated on both VLM backbones under consistent settings, where we fix the decoding temperature at 0 and use only the base prompts provided by each dataset to ensure fair comparisons. Since REVERSE does stochastic sampling at inference time, we report the mean performance over 100 bootstrapped runs for robustness. The exact numbers with 95% confidence intervals are provided in Appendix E.
Our evaluation dataset covers several standard VQA tasks aimed at assessing visual hallucination, with a primary focus on image captioning and open-ended question answering. While discriminative tasks, binary (Yes/No) questions targeting object, attributes, and spatial understanding, are also common, backtracking provides limited benefit in such settings, which have been noted to offer less diagnostic insight into VLM hallucinations [7, 39]. We include results on these tasks in Appendix E for completeness.
For image captioning, we use CHAIR-MSCOCO [38, 53] and the generative subset of AMBER [45] (denoted as AMBER-G). CHAIR-MSCOCO evaluates object hallucination using the CHAIR score, which measures the degree of misalignment between objects mentioned in a model-generated caption and objects actually present in the image. It is defined as 1 minus the intersection over union (IoU) between the sets of mentioned and ground-truth objects. We report both CHAIRi, which aggregates the CHAIR score across all object instances, and CHAIRs, which quantifies the proportion of images where at least one hallucination occurs. For AMBER-G, we report four key metrics: CHAIR, Coverage (Cover), Hallucination (Hall), and Cognition (Cog). CHAIR is the same as CHAIRi above, and Coverage measures how well the caption mentions all objects in the image, similar to recall. The definitions of the remaining two metrics and further details on these datasets are provided in Appendix D.
For open-ended question answering, we evaluate on MMHal-Bench [40] and HaloQuest [47], which generally test VLMs on false-premise questions, questions with insufficient visual evidence, and visually complex queries. Following standard evaluation protocols, we assess MMHal-Bench responses using gpt-4-0314 and HaloQuest with Gemini-1.5-Pro, as the original paper used Gemini-1.0-Pro, which is no longer available. These benchmarks require models to generate free-form text answers, testing their ability to comprehend and reason about visual content in an open-ended manner.
| Method | Avg. Acc. | FP Acc. | VC Acc. | IC Acc. |
| LLaVA-v1.5 7B | ||||
| None† | 22.6 | 17.1 | 39.5 | 10.7 |
| DoLA† [15] | 22.9 | 17.2 | 40.1 | 11.6 |
| HALVA† [39] | 23.9 | 21.1 | 37.4 | 10.7 |
| REVERSE() | 30.7 | 31.8 | 31.5 | 26.9 |
| REVERSE() | 32.3 | 29.4 | 18.7 | 58.8 |
| LLaVA-MORE 8B | ||||
| None† | 22.4 | 15.8 | 43.4 | 7.4 |
| DoLA† [15] | 22.8 | 15.5 | 45.1 | 7.4 |
| REVERSE() | 26.7 | 30.0 | 31.3 | 11.7 |
| REVERSE() | 36.7 | 39.5 | 30.9 | 38.1 |
| Qwen2.5-VLFT 3B | ||||
| None† | 33.5 | 25.4 | 51.6 | 26.4 |
| DoLA† [15] | 27.4 | 16.5 | 51.1 | 19.0 |
| REVERSE() | 45.1 | 42.9 | 41.8 | 55.5 |
| GPT-4o | 63.2 | 65.2 | 55.2 | 68.7 |
| Gemini 1.5 Pro | 77.9 | 83.7 | 56.3 | 92.5 |
| Base VLM | Method | Score | Hall. Rate |
| LLaVA-v1.0 7B | LLaVA-RLHF [40] | 2.05 | 0.68 |
| LLaVA-v1.5 7B | None [34] | 2.11 | 0.54 |
| HACL [23] | 2.13 | 0.50 | |
| HA-DPO [54] | 1.97 | 0.60 | |
| EOS [53] | 2.03 | 0.59 | |
| HALVA [39] | 2.25 | 0.54 | |
| DoLA† [15] | 2.33 | 0.56 | |
| Woodpecker† [51] | 2.19 | 0.58 | |
| REVERSE() | 2.56 | 0.47 | |
| REVERSE() | 3.28 | 0.30 | |
| LLaVA-MORE 8B | None† | 2.50 | 0.53 |
| DoLA† [15] | 2.54 | 0.51 | |
| Woodpecker† [51] | 2.28 | 0.58 | |
| REVERSE() | 2.28 | 0.54 | |
| REVERSE() | 2.93 | 0.40 | |
| Qwen2.5-VLFT 3B | None† | 2.89 | 0.43 |
| DoLA† [15] | 2.72 | 0.46 | |
| REVERSE() | 3.15 | 0.29 |
4.1 Experimental Results
Image Captioning Tasks Table 1 presents results on image captioning tasks. With the default parameters, our method achieves the best results on LLaVA-series and Qwen2.5-VL, reducing the CHAIRi value by up to 12% on CHAIR-MSCOCO and AMBER-G compared to the best existing methods. On LLaVA-v1.5-7B, training-free methods underperform relative to fine-tuned approaches. DoLA [15] reduces hallucinations for LLaMA-based models but does not generalize well to Qwen2.5-VL. EOS [53] performs well on AMBER-G, likely because it encourages the model to produce more concise captions, leading to less informative outputs and reduced coverage. The post-hoc refinement method, Woodpecker [51], employs a multi-stage process for verification and correction; however, its one-time correction approach may suffer from error propagation, limiting its effectiveness. In contrast, REVERSE generalizes across models and supports threshold-based tuning to balance hallucination and coverage (more studies in subsection 4.2). Figure 4 presents qualitative results using four different methods. LLaVA-v1.5-7B, OPERA, and Woodpecker hallucinate non-existing objects, while REVERSE can generate the correct caption without reducing caption length too much.
Open-ended Question Answering We evaluate REVERSE on MMHal-Bench and HaloQuest, two open-ended VQA benchmarks containing lots of questions with false premises or insufficient context. These examples require the model to either refuse or correct the query. For these questions, we observed that REVERSE often produces empty responses and we interpret this behavior as the model identifying the query as unanswerable. In all such cases, we apply query rewriting with the prompt: “For this question, please point out the false premises or note what information is missing, rather than answering it directly.” The complete mechanism for handling unanswerable questions is given in Appendix F.
As shown in Table 3 and Table 2, REVERSE improves accuracy by up to 10% on MMHal-Bench and 34% on HaloQuest compared with the SOTA models using default hyperparameters ( for LLaVA series and for Qwen2.5-VL). Most of the gains come from better handling of false-premise and insufficient-context questions. However, performance on visually challenging questions decreases as the model adopts a more cautious approach and avoids speculative answers, even when they may be correct. Additional experiments with LLaVA show that lowering the threshold further (e.g., ) increases conservativeness and boosts performance on ambiguous queries, without requiring task-specific fine-tuning.
4.2 Discussions
We conduct additional experiments on two LLaVA models to examine key aspects of our method, including ablation studies, trade-offs between expressiveness and performance, efficiency versus accuracy, and the effect of temperature. Potential limitations and broader social impacts are discussed in Appendix C.
Ablation Studies We conduct ablation studies to evaluate the contributions of different components of our method, as shown in Table 4. Comparing the first and second rows, we observe that hallucination-aware training alone already improves performance across all metrics, outperforming existing VLMs. We hypothesize that this improvement arises from the model’s ability to contrast positive and negative phrases, effectively learning to distinguish between </CN> and </UN> during training—a mechanism that may be similar to DPO [37]. This finding suggests a potential research direction for future work. Interestingly, even a naive rejection sampling strategy reduces CHAIR hallucination scores by 1.2. When combined with query rewriting, the coverage improves by a further 1.2 points, indicating that rewriting helps the model explore alternative phrasing and correct itself more effectively.
Trade-offs Between Expressiveness and De-Hallucination As discussed in subsection 3.3, a key component of our retrospective resampling method is the predefined threshold . When the predicted probability of </UN> exceeds , the model triggers backtracking and self-correction. Figure 5 presents an analysis of the effect of on two VLMs. The 2D plot illustrates the performance trade-off between CHAIR (hallucination metric) and coverage across different threshold values. Based on our experiments, we selected as a global threshold for image captioning and other generative tasks, as it represents the peak of the performance frontier. Compared with prior hallucination reduction methods, REVERSE enables us to dynamically adjust the balance between expressiveness and trustworthiness. Notably, with a relatively high threshold of , our method already surpasses the base VLMs (LLaVA-v1.5 and LLaVA-MORE) in both hallucination reduction and content coverage. Moreover, with a lower threshold (), our model can even outperform GPT-4V in hallucination control.
| Components | CHAIR | Cover | Hall | Cog |
| LLaVA-v1.5-7B | 7.8 | 51.0 | 36.4 | 4.2 |
| + Hall-aware Training | 7.2 | 53.2 | 36.3 | 3.4 |
| + Rejection Sampling | 6.0 | 51.0 | 30.5 | 3.0 |
| + Query Rewriting | 6.0 | 52.2 | 30.4 | 3.0 |
Trade-offs Between Inference Efficiency and De-Hallucination REVERSE combines online self-verification with re-generation, introducing minimal computational overhead. To assess this, we analyze 1,004 samples from the AMBER-G test set. Verification is triggered in only 63% of cases and is highly efficient, relying on lightweight logit-based checks as opposed to a heavy pipeline in prior post-hoc verification method [51]. In the remaining 37% where verification is skipped, the CHAIR score is already low (5.1 vs. the average of 7.2), indicating that our verification can identify and skip high-quality outputs.
Among the 63% of cases where backtracking is triggered, no grounded noun phrases are mistakenly corrected to hallucinated ones. Approximately 1.5% of all cases contain hallucinations, over half of which are resolved with a single re-generation round. Furthermore, re-generation can be optimized by reusing KV-cache [43, 25], eliminating the need to recompute the full prefix. Overall, the self-correction loop in REVERSE is computationally affordable, and its effectiveness is demonstrated in Table 1, Table 2, and Table 3.
Impact of Temperature on Hallucination and Coverage We also analyze how temperature settings affect hallucination rates and object coverage in generated outputs. In this experiment, we use LLaVA-v1.5-7B as the backbone and increase the number of local correction attempts () while keeping all other parameters in REVERSE unchanged. As shown in Figure 6, REVERSE is robust to increasing temperature. For tasks such as image captioning, a higher temperature is often desirable to enhance diversity in generated descriptions. However, existing methods not only suffer from increased hallucinations at higher temperatures but also exhibit a decline in object coverage. In contrast, REVERSE balances expressiveness and reliability—slightly reducing hallucination while improving object coverage as temperature increases.
5 Conclusion
In this paper, we introduced REVERSE, a framework that reduces hallucinations in Vision-Language Models by combining hallucination-aware training with retrospective resampling. REVERSE achieves up to 12% improvement on image CHAIR-MSCOCO and 34% on HaloQuest over existing SOTA methods. While preliminary, REVERSE highlights the potential of self-correction in multimodal models. Future work may explore integrating structured verification and causal reasoning to further reduce hallucinations. As multimodal AI progresses, we see self-verification and retrospective techniques as a promising direction for building more trustworthy systems.
Acknowledgments
We thank Shalini Ghosh and Konpat Preechakul for their invaluable feedback during the discussions. Authors, as part of their affiliation with UC Berkeley, were supported in part by the National Science Foundation, US Department of Defense, and/or the Berkeley Artificial Intelligence Research (BAIR) industrial alliance program, as well as gifts from Amazon. Sky Computing Lab is supported by gifts from Accenture, AMD, Anyscale, Cisco, Google, IBM, Intel, Intesa Sanpaolo, Lambda, Microsoft, NVIDIA, Samsung SDS, SAP, and VMware. This research was also developed with funding from the Defense Advanced Research Projects Agency (DARPA) under Contract No(s). FA8650-23-C-7316 and HR0011-25-3-0133. The views, opinions and/or findings expressed are those of the author and should not be interpreted as representing the official views or policies of any sponsor, the Department of Defense, or the U.S. Government.
Appendix
The appendix consists of the following further discussion:
Appendix A provides links to the released code, model checkpoints, and dataset.
Appendix B describes the dataset that we constructed containing (</UN>) and (</CN>) tokens.
Appendix C discusses the potential limitations and social impact of this work.
Appendix D describes the evaluation datasets and metrics.
Appendix E provides more qualitative and quantitative results.
Appendix F details the training and evaluation implementation for REVERSE.
Appendix A Code and Model Release
The project website is available at: https://reverse-vlm.github.io. The code for REVERSE is released under the MIT license at https://github.com/tsunghan-wu/reverse_vlm. It builds upon the Apache 2.0-licensed codebases of LLaVA [33] and LLaVA-MORE [16], as well as the Qwen license associated with Qwen2.5-VL [6].
We also release model checkpoints of REVERSE (based on LLaVA-v1.5, LLaVA-MORE, and Qwen2.5-VL), along with a 1.3M-sample semi-synthetic dataset, at Hugging Face. Both the checkpoints and dataset are released under the MIT license. The dataset contains elements adapted from LLaVA [33], which is licensed under the Creative Commons Attribution 4.0 International License. Use of the dataset complies with OpenAI’s usage policy.
Appendix B Dataset Details
In the construction of the dataset, the automatically annotated phrases include all noun phrases (multilingual support) and a set of predefined keywords relevant to the task, such as “Yes” or “No.” We maintain a skip list to prevent the model from marking common but uninformative phrases, like “in the image.” Noun phrases with their corresponding prefix prepositions are automatically extracted using Part-of-Speech (POS) tagging tools [21], and the full data curation pipeline is in Figure A.1. The skip list is in Figure A.3. In this work, we only annotate model response phrases rather than the questions. We also show the prompt for constructing negative phrases for general questions in the following pages.
To further enhance dataset quality and balance, we validate that all tokens are tagged properly and then redistribute positive and negative answers. The final image distribution of our data composition is the same as the original LLaVA-v1.5-665k dataset as shown in Figure A.2.
Appendix C Limitations and Societal Impact
Our 1.3M Dataset:
We synthesize a 1.3M-sample, hallucination-aware instruction tuning dataset. While effective as a proof-of-concept (shown in Table 1, Table 2), it has several limitations. First, as with LLaVA’s dataset, ours lacks coverage of edge cases (e.g., insufficient context, false premises) and noisy labels, which is an issue recently addressed in newer datasets [13, 41]. Second, our data augmentation uses GPT-4o-mini, which may introduce bias or limited coverage. Finally, our dataset includes subsets from existing sources like MS-COCO, which contain known biases (e.g., gender, race, geography) [19, 9, 20, 44]. Future work should aim to develop more comprehensive, higher-quality datasets with reduced bias.
REVERSE-VLMs:
REVERSE is a generalizable framework applicable to models like LLaVA-v1.5, LLaVA-MORE, and Qwen2.5-VL. It significantly reduces hallucination with minimal loss in expressiveness and efficiency (shown in Table 1, Table 2). However, it does not improve performance on discriminative VQA tasks (see Appendix D) as the backtracking methods generally didn’t help further reasoning. Future work can explore integrating REVERSE with existing de-hallucination techniques to improve such tasks. As a well studied problem, REVERSE carries risks of misuse just like other VLMs [8, 33, 34]. We adopt existing safety mechanisms from upstream models (i.e., LLaVA-v1.5 [34], LLaVA-MORE [16], and Qwen2.5-VL [6]), and provide a fully open-source release with training details.
what, where, which, who, whom, whose, why, how,
What, Where, Which, Who, Whom, Whose, Why, How,
that, this, these, those, That, This, These, Those,
he, she, it, we, you, they, me, him, her, us, them, I,
He, She, It, We, You, They, Me, Him, Her, Us, Them, I,
my, your, his, her, its, our, their, mine, yours, ours, theirs,
My, Your, His, Her, Its, Our, Their, Mine, Yours, Ours, Theirs,
a, an, the, A, An, The,
in the image, the image, The image, In the image,
in the picture, the picture, The picture, In the picture.
Appendix D Evaluation Datasets & Metrics
To evaluate how REVERSE reduces visual hallucination through backtracking, we use two image captioning benchmarks and two open-ended VQA datasets. While not the main focus of this paper, we also report results on standard discriminative VQA benchmarks commonly used for hallucination evaluation. Below, we describe each of these benchmark datasets in detail.
CHAIR-MSCOCO:
MS COCO (Microsoft Common Objects in Context) [30] is a large-scale dataset designed for object detection, segmentation, and captioning tasks in computer vision. It contains over 330,000 images, with more than 200,000 labeled images spanning 80 object categories. The dataset includes detailed instance annotations, allowing for precise object localization and segmentation. Additionally, MS COCO provides five human-generated captions per image, making it a popular benchmark for image captioning and vision-language models (VLMs).
The CHAIR-MSCOCO benchmark was first introduced by Rohrbach et al. [38], which uses the full MSCOCO validation set to evaluate hallucination in vision-language models using the CHAIR score. In this work, we follow the evaluation protocol of Yue et al. [53] and assess a subset of 500 captions for efficient benchmarking.
AMBER:
AMBER (An LLM-free Multi-dimensional Benchmark for MLLMs Hallucination Evaluation) [45] is a comprehensive evaluation framework designed to assess hallucination phenomena in Multi-modal Large Language Models (MLLMs). Unlike previous benchmarks that often rely on human or advanced LLM evaluations, AMBER offers an automated, low-cost approach to evaluate both generative and discriminative tasks. It addresses three key dimensions of hallucination: existence, attribute, and relation.
In this work, we refer to the generative subset as AMBER-G and the discriminative subset as AMBER-D. The AMBER-D subset comprises binary Yes/No questions and is evaluated using the F1 score, consistent with the POPE benchmark. Evaluation on AMBER-G is more complex and involves multiple metrics, as described below.
CHAIR [38] measures the percentage of hallucinated objects in a scene, and is defined as:
where ={, , , } is an annotated list of objects, and ={, , , } are nouns extracted from the captions using NLTK.
Cover measures the object coverage of responses, namely, the proportion of objects mentioned in the response relative to the objects identified in the , and is defined by:
Hal is a more general metric, measuring the portion of responses containing hallucinations (similar to CHAIRs). It is defined as:
Cog measures how similar the hallucinations that a model generates are to human hallucinations. It is defined as:
for a set of target hallucinatory objects ={, , , }.
HaloQuest:
HaloQuest [47] is a VQA dataset designed to evaluate and mitigate hallucination in VLMs. The evaluation set consists of over 600 examples, featuring both real images from the OpenImages dataset and synthetic images generated using tools like Midjourney. The dataset focuses on three categories of questions: those with false premises, those lacking sufficient context, and visually challenging ones, aiming to trigger common hallucination scenarios in VLMs. Performance on Haloquest is measured using “accuracy” as evaluated by Gemini-1.0 Pro; however since this model is no longer available, we leverage Gemini 1.5-Pro in this paper.
MMHal-Bench:
MMHal-Bench [40] is an evaluation benchmark specifically designed to assess hallucination phenomena in Large Multimodal Models (LMMs). It comprises 96 challenging image-question pairs sourced from the OpenImages dataset, each accompanied by corresponding ground-truth answers and detailed image content annotations. The benchmark focuses on penalizing hallucinations by evaluating model responses against these ground truths. Performance on MMHal-Bench is measured using “score” (ranging from 0-6) and “hallucination rate” (ranging from 0-1) as evaluated by a GPT model.
| Base VLM | Method | AMBER-D | POPE | MME-Hall |
| LLaVA-v1.5 7B [34] | None | 74.7 | 85.9 | 648.3 |
| VCD [27] | - | 84.5 | 604.7 | |
| EOS [53] | 75.6 | 86.0 | 606.7 | |
| OPERA† [22] | 74.8 | 85.5 | 592.3 | |
| DoLA [15] | 74.5 | 85.7 | 656.7 | |
| HA-DPO [54] | 78.1 | 86.9 | 618.3 | |
| MEMVR [56] | - | 85.9 | 648.3 | |
| AGLA [3] | - | 86.0 | 640.0 | |
| HALVA [39] | 83.4 | 84.8 | 665.0 | |
| Woodpecker [51] | 67.0 | - | 366.7 | |
| REVERSE() | 74.2 | 85.9 | 601.6 | |
| LLaVA-MORE 8B [16] | None | 71.6 | 85.1 | 678.3 |
| DoLA [15] | 72.0 | 85.2 | 683.3 | |
| REVERSE() | 69.3 | 84.4 | 657.6 | |
| Qwen2.5-VLFT 3B [6] | None | 87.7 | 87.1 | 550.4 |
| DoLA [15] | 89.0 | 78.8 | 555.6 | |
| REVERSE() | 85.7 | 86.5 | 589.5 |
POPE:
The Polling-based Object Probing Evaluation (POPE) [29] is a benchmark designed to assess object hallucination in vision-language models (VLMs). Unlike traditional instruction-based evaluations, POPE employs a polling-based query method, prompting LVLMs with simple yes-or-no questions about the presence of specific objects in images. This approach converts the evaluation into a binary classification task, allowing for more stable and flexible assessment of object hallucination. Performance on POPE is measured in F1 score following LLaVA’s standard [33].
MME-Hall:
MME-Hall [18] is a specialized subset of the Multimodal Large Language Model Evaluation (MME) benchmark, focusing specifically on assessing object-related hallucinations in multimodal large language models (MLLMs). It evaluates models across four key dimensions: object existence, counting, positional accuracy, and color recognition. Most questions in this subset require binary Yes/No answers or brief responses in short phrases.
Appendix E Additional Results
Discriminative Tasks:
We report results on discriminative hallucination benchmarks including AMBER-D, POPE, and MME-Hall in Appendix D. These benchmarks consist of binary classification tasks such as Yes or No questions, where the impact of retrospective resampling is naturally limited. Since the answer space is minimal, often restricted to a single token, rethinking the output using a clarification hint like “(potential incorrect phrases Yes/No)” described in subsection 3.3 generally does not provide meaningful improvement.
In this setting, we set the hallucination detection threshold to for computational efficiency. Increasing the threshold beyond 0.5 does not significantly affect the results. Across all datasets, REVERSE performs comparably to existing baselines but does not show substantial gains. This is likely due to the binary nature of these tasks, which offer limited opportunity for backtracking or resampling to influence the outcome.
By contrast, for open-ended and captioning tasks, we adopt lower thresholds (0.003 for LLaVA-based models and 0.01 for Qwen2.5-VL) to account for the much larger answer space. For instance, in a sentence such as “There is a ,” the blank can be filled with many possible words, and both non-hallucinatory tokens generally have relatively low probabilities. In such cases, setting an appropriately low threshold is critical for effective hallucination detection. As discussed in section 4, we use a fixed threshold per model across datasets to ensure a fair comparison. We further reflect on its effectiveness in the discussion subsection. We leave further exploration of threshold calibration and reasoning strategies for binary tasks to future work.
Bootstrapped Evaluation Results:
As described in section 4, we apply 100-round bootstrapping to account for variability introduced by sampling during inference, particularly for REVERSE and smaller datasets such as MM-Hal, which contains only 96 samples. Results are reported in Table A.2, Table A.3, and Table A.4.
Across all captioning and open-ended VQA tasks, REVERSE consistently demonstrates strong robustness and significantly outperforms the base VLM, with margins exceeding the 95% confidence interval. For discriminative tasks, the performance of REVERSE mostly remains within the 95% confidence interval of the base model, indicating comparable performance.
Qualitative Examples:
Figure A.4 presents additional qualitative comparisons between REVERSE and prior methods. Words highlighted in red indicate hallucinations. Similar to Figure 4, REVERSE significantly reduces hallucinations. In particular, the fourth row clearly illustrates that when the model is uncertain, it avoids adding speculative or unsupported content.
| Base VLM | Method | CHAIR-MSCOCO | AMBER-G | ||||
| CHAIR | Cover | Hall | Cog | ||||
| LLaVA-v1.5 7B [34] | Base VLM | 15.4 | 50.0 | 7.8 | 51.0 | 36.4 | 4.2 |
| REVERSE() | 10.3(8.94–11.68) | 37.0(32.90–41.50) | 6.0(5.4–6.5) | 52.2(51.1–53.5) | 30.4(27.8–32.9) | 3.0(2.5–3.4) | |
| REVERSE() | 6.1(4.53–7.60) | 13.6(10.80–16.71) | 4.0(2.3–5.9) | 26.9(24.1–30.6) | 10.2(6.6–14.2) | 0.9(0.4–1.5) | |
| LLaVA-MORE 8B [16] | Base VLM | 14.4 | 52.0 | 7.8 | 53.1 | 36.6 | 3.9 |
| REVERSE() | 12.2(10.55–13.81) | 42.4(38.09–46.02) | 6.5(5.9–7.1) | 54.8(53.6–56.2) | 35.5(32.4–38.9) | 3.9(3.3–4.4) | |
| REVERSE() | 8.4(7.16–10.06) | 25.2(21.39–28.60) | 5.1(4.5–5.6) | 38.9(37.3–40.4) | 20.8(18.6–23.2) | 2.1(1.7–2.5) | |
| Qwen2.5-VLFT [6] | Base VLM | 12.2 | 45.8 | 7.7 | 51.7 | 35.9 | 4.1 |
| REVERSE() | 10.5(9.22–11.90) | 39.4(35.40–43.91) | 7.5(6.6–8.1) | 51.5(50.2–52.8) | 34.4(31.5–37.2) | 3.6(3.1–4.2) | |
| Backbone | Method | MMHal | HaloQuest | ||||
| Score | Hall. Rate | Avg Acc. | FP Acc. | VC Acc. | IC Acc. | ||
| LLaVA-v1.5 7B | Base VLM | 2.11 | 0.54 | 22.6 | 17.1 | 39.5 | 10.7 |
| REVERSE() | 2.56(2.22–3.01) | 0.47(0.35–0.56) | 30.7(27.4–35.2) | 31.8(27.5–35.7) | 31.5(25.2–38.1) | 26.9(20.8–34.7) | |
| REVERSE() | 3.28(2.86–3.72) | 0.30(0.20–0.40) | 32.3(29.4–36.5) | 29.4(25.1–35.7) | 18.7(13.5–24.6) | 58.8(50.0–67.6) | |
| LLaVA-MORE 8B | Base VLM | 2.50 | 0.53 | 22.4 | 15.8 | 43.4 | 7.4 |
| REVERSE() | 2.28(1.96–2.68) | 0.54(0.44–0.62) | 26.7(23.4–30.0) | 30.0(26.4–35.2) | 31.3(25.3–38.2) | 11.7(6.1–16.0) | |
| REVERSE() | 2.93(2.53–3.22) | 0.40(0.31–0.51) | 36.7(33.9–39.8) | 39.5(34.8–45.0) | 30.9(26.6–37.2) | 38.1(31.6–46.9) | |
| Qwen2.5-VLFT 3B | Base VLM | 2.89 | 0.43 | 33.5 | 25.4 | 51.6 | 26.4 |
| REVERSE() | 3.15(2.80–3.45) | 0.29(0.21–0.40) | 45.1(40.9–48.9) | 42.9(36.5–48.6) | 41.8(34.8–49.1) | 55.5(44.6–65.0) | |
| Backbone | Method | AMBER-D | POPE | MME-Hall |
| LLaVA-v1.5 7B | Base VLM | 74.7 | 85.9 | 648.3 |
| REVERSE | 74.2(73.4–75.2) | 85.9(82.8–88.5) | 601.60(555.00–642.54) | |
| LLaVA-MORE 8B | Base VLM | 71.6 | 85.1 | 678.3 |
| REVERSE | 69.3(68.5–70.3) | 84.4(81.8–86.9) | 657.63(615.71–696.83) | |
| Qwen2.5-VLFT 3B | Base VLM | 85.0 | 87.1 | 550.4 |
| REVERSE | 85.7(85.1–86.1) | 86.5(84.3–88.8) | 589.48(544.92–635.00) |
Appendix F Implementation Details
F.1 Training Details
For both LLaVA-v1.5-7B and LLaVA-MORE, we initialize from pretrained language models (Vicuna-1.5-7B and Llama-3.1-8B-Instruct, respectively), along with their corresponding visual projectors and CLIP-ViT-L/14-336 vision encoders. Following the standard LLaVA setup, we perform instruction fine-tuning using our 1.3M multi-image dataset for one epoch with LoRA (rank = 128, ). We adopt the modified cross-entropy loss defined in subsection 3.2 and train using the AdamW optimizer. The learning rate is set to 2e-5 for the visual projector, and (2e-4, 1e-4) for the LoRA parameters of LLaVA-v1.5-7B and LLaVA-MORE, respectively. The CLIP backbone is kept frozen. We use a global batch size of 128 with no gradient accumulation. Training takes 24 hours for LLaVA-v1.5-7B and 36 hours for LLaVA-MORE on 8× A100 80GB GPUs using DeepSpeed ZeRO-2.
For Qwen2.5-VL, since the instruction tuning dataset is unavailable, we finetune the released model directly. To enable apples-to-apples comparison, we apply our REVERSE finetuning on the same 100k subset used for LLaVA-FT. Unlike the LLaVA variants, we finetune the full 3B Qwen2.5-VL model (without LoRA) using the same modified cross-entropy loss. We use the AdamW optimizer with a learning rate of 5e-5, freeze the CLIP encoder, and set the batch size to 128 with no gradient accumulation. Training takes 3 hours on 4× A100 80GB GPUs using DeepSpeed ZeRO-3.
F.2 Full Decoding Algorithm
The full decoding algorithm used in REVERSE is shown in Algorithm 1. For captioning and discriminative tasks, we directly apply this algorithm.
For open-ended tasks such as MMHal-Bench and HaloQuest, as described in section 4, we adopt a two-stage decoding process. In the first round, REVERSE runs the standard retrospective resampling algorithm. Since many of these queries contain false premises or lack sufficient information, the model is expected to abstain from answering and return a blank response. In such cases, we initiate a second round of inference with the query modified as:
Q := Q + "For this question, please point out the false premises or note what information is missing, rather than answering it directly."
This prompting strategy requires no additional training and enables the model to handle underspecified or invalid queries more effectively.